# Bayesian Perspective for Orientation Determination in Cryo-EM with Application to Structural Heterogeneity Analysis

### v2

@@ 4.2 Empirical results for volume reconstruction and the “Einstein from Noise” phenomenon
Figure 4: Comparison of 2D image recovery using the MMSE and the MLE rotation estimators. Iterative image recovery procedures with the MMSE estimator ($\hat{g}_{\text{MMSE}}$) and MLE estimator ($\hat{g}_{\text{MLE}}$) are defined in (4.4) and (4.6), respectively. The experiment employs a template image of Einstein and a ground truth image of Newton, both rotated in 2D over a uniform polar grid with $L=30$ samples. The reconstructed images within the dark-orange rectangle (right panel) show superior performance with the MMSE rotation estimator, with Pearson cross-correlation (PCC) provided for each reconstructed image. The MLE and MMSE reconstructions are nearly identical at high SNR ($\sigma\to 0$), as predicted by Proposition 2.2. The SNR values used for the panels (from right to left) are $10^{-2}$, $4\times 10^{-3}$, $2\times 10^{-3}$, $7\times 10^{-4}$, and $2\times 10^{-4}$. At very low SNR ($\sigma\to\infty$), the “Einstein from Noise” effect appears, where the estimator resembles the template image of Einstein rather than the underlying truth of Newton. In intermediate SNR ranges, using the MMSE estimator in the iterative step clearly outperforms the MLE estimator.
Figure 5: Comparison of 3D structure reconstruction in cryo-ET subtomograms averaging using the MMSE and MLE rotation estimators. Iterative structure reconstruction procedures with the MMSE estimator ($\hat{g}_{\text{MMSE}}$) and MLE estimator ($\hat{g}_{\text{MLE}}$) are defined in (4.4) and (4.6), respectively. At high SNR levels, a low-resolution structure emerges due to finite grid sampling of the rotation group $\mathsf{SO}(3)$, effectively acting as a low-pass filter. The 3D reconstruction using the MMSE estimator consistently outperforms the reconstruction using the MLE estimator. For high SNR conditions (i.e., $\sigma\to 0$), both estimators yield similar 3D structures, as expected from Proposition 2.2. The SNR values used for the panels (from right to left) are $10^{-2}$, $2\times 10^{-3}$, $7\times 10^{-4}$, and $2\times 10^{-4}$, with a volume size of $32\times 32\times 32$. The boxes highlighted in purple resemble the true input structure (Ribosome-S80 wong2014cryo ), while the orange-highlighted boxes are more similar to the initial template (beta-galactosidase bartesaghi2014structure ), illustrating the “Einstein from Noise” phenomenon. Notably, at very low SNRs, the “Einstein from Noise” effect is evident with the MLE estimator but not with the MMSE estimator.
@@
A few observations can be made from Figures 4 and 5. First, in the case of high SNR (i.e., as $\sigma\to 0$), the volume reconstruction is similar whether using the MLE estimator or the MMSE estimator. This similarity is theoretically supported by Proposition 2.2. However, as the SNR decreases, the reconstructions diverge, with the volumes reconstructed using the MMSE estimator showing a better correlation with the true volume. Second, in scenarios of extremely low SNR, where the structural signal is nearly nonexistent, the phenomenon known as “Einstein from Noise” manifests in both 2D and 3D contexts. This phenomenon pertains to the inherent model bias within the reconstruction procedure, specifically in relation to the initial templates. In such cases, the reconstructed volume exhibits structural similarities to the initial template, even though the observations do not substantiate this outcome. The generation of a structured image from entirely noisy data has attracted considerable attention, particularly during a significant scientific debate regarding the structure of an HIV molecule mao2013molecular ; henderson2013avoiding ; van2013finding ; subramaniam2013structure ; mao2013reply ; for a comprehensive description and statistical analysis, see balanov2024einstein ; balanov2024confirmation ; balanov2025expectation ; balanov2025structure . Notably, our empirical evidence suggests that the “Einstein from Noise” phenomenon is more pronounced when adopting the MLE estimator compared to the MMSE estimator, implying that the MMSE approach is less vulnerable to the choice of the initial template. Furthermore, our experiment suggests that the advantage of using the Bayesian MMSE estimator over the MLE estimator is more significant in 3D structure reconstruction tasks compared to 2D image recovery, where the 3D setting is a problem of greater interest to researchers in structural biology.
@@ 5 Structural heterogeneity analysis
Understanding heterogeneity is central to revealing the dynamic behavior of macromolecular complexes in cryo-EM. Unlike traditional 3D classification methods that assume a small number of discrete states, modern approaches aim to capture continuous structural variability by embedding projection images into a latent conformational space. Here, we adopt the RECOVAR framework gilles2025cryo , which performs heterogeneity analysis using principal component analysis (PCA) based on a regularized estimate of the conformational covariance matrix. The key idea is to infer the covariance of 3D volumes directly from noisy 2D projection images, leveraging the fact that although the volumes themselves are unobserved, their projections contain sufficient statistical information.
@@ 5.1 Fixed-pose methods
A central computational assumption in RECOVAR is that the poses $\{\phi_{i}\}$ of all particle images are known and fixed, typically obtained from a consensus refinement procedure. This “fixed-pose” assumption underlies many modern heterogeneity analysis pipelines, including 3D Classification in cryoSPARC punjani2017cryosparc , 3D Variability Analysis (3DVA) punjani20213d , 3DFlex punjani20233dflex , cryoDRGN zhong2021cryodrgn , CryoDRGN-AI-fixed levy2024end , Opus-DSD luo2023opus , and RECOVAR gilles2025cryo . These methods are collectively referred to as fixed-pose methods, since they treat poses as known inputs during downstream inference. Among them, RECOVAR has recently emerged as one of the most effective approaches. According to the CryoBench evaluation jeon2024cryobench , which benchmarks multiple methods across diverse synthetic and experimental datasets, RECOVAR consistently achieves state-of-the-art performance in both structural resolution and latent space recovery. This motivates our choice to build upon the RECOVAR framework.
@@ 5.2 The mathematical model
The forward model (gilles2025cryo, , Eq. (1)) used in RECOVAR is expressed in the Fourier domain as
@@
where $\tilde{y}_{i}\in\mathbb{C}^{N^{2}}$ is the observed Fourier-transformed image (with $N$ denoting the grid size along each spatial axis), $\tilde{V}_{i}\in\mathbb{C}^{N^{3}}$ is the (unknown) 3D Fourier volume corresponding to a particular conformation, $\hat{P}(\phi_{i})$ is the tomographic projection operator from 3D to 2D after a rigid-body motion $\phi_{i}=(g_{i},t_{i})$ with $g_{i}\in\mathrm{SO}(3)$ a rotation and $t_{i}\in\mathbb{R}^{2}$ an in-plane shift, $C_{i}$ is the contrast transfer function (CTF), and $\tilde{\varepsilon}_{i}$ is additive noise.
@@
where the expectation is taken over the conformational distribution. The objective of RECOVAR is to estimate both $\mu$ and $\Sigma$ from the observed projections $\tilde{y}_{i}$} with given poses $\phi_{i}$}, using regularized least-squares minimization (gilles2025cryo, , Eqns. (2)–(3)).
@@ 5.3 Numerical experiments
In our experiments, we consider a synthetic one-dimensional conformational transition, a standard benchmark for continuous heterogeneity analysis methods such as cryoDRGN zhong2021cryodrgn and RECOVAR gilles2025cryo . The corresponding ground-truth density maps along the conformational coordinate are shown in Figure 6(a). To generate the projection images, the underlying conformational states were uniformly sampled along the one-dimensional transition, and molecular orientations were sampled from a uniform distribution over $\mathrm{SO}(3)$. Figure 6(b) further displays a representative noisy projection under high-noise conditions.
To avoid storing and manipulating the full high-dimensional covariance matrix $\Sigma\in\mathbb{C}^{N^{3}\times N^{3}}$, RECOVAR first estimates a low-dimensional subspace of rank $r\ll N^{3}$ in which the covariance is approximated. Following (gilles2025cryo, , Appendix A.2), we estimate $\hat{\Sigma}$ via regularized least-squares, then form a rank-$d$ approximation $\hat{\Sigma}_{\mathrm{col}}$ by selecting a subset of columns using a greedy SNR-based criterion that ensures both high SNR and low inter-column correlation. An orthonormal basis $\tilde{U}\in\mathbb{C}^{N^{3}\times r}$ spanning $\hat{\Sigma}_{\mathrm{col}}$ is obtained via randomized SVD. The conformational mean $\hat{\mu}$ is also estimated at this stage. Each conformational volume is then represented as
@@
where $\hat{\mu}$ is the estimated conformational mean from the first stage, and $z_{i}\in\mathbb{C}^{r}$ are low-dimensional coordinates. In RECOVAR, these coordinates are not explicitly recovered; their covariance matrix $\hat{\Sigma}_{\tilde{U}}\in\mathbb{C}^{r\times r}$ is estimated directly via a reduced least-squares problem (Eq. (14) in gilles2025cryo ), yielding a low-rank covariance approximation that characterizes the variability of $z_{i}$. The eigen-decomposition of $\hat{\Sigma}_{\tilde{U}}$ then provides the estimated principal components and eigenvalues in the reduced space. The complete procedure is detailed in (gilles2025cryo, , Algorithm 1).
We perform our analyses on a controlled synthetic dataset of 30,00030{,}000 projection images ($N=128$), simulated from a one-dimensional conformational transition discretized into $50$ equally spaced states to approximate continuous heterogeneity gilles2025cryo . We first evaluate the effect of orientation estimation on covariance recovery. The study is conducted in the high-noise regime of RECOVAR (Figure 6(b)), where even with ground-truth poses, the top $30$ principal components capture only about $40\%$ of the variance due to finite-sample and high-noise effects. We set $r=50$ and compare three pose sources: ground truth, MMSE estimates, and MLE estimates, keeping all other steps identical. For fixed-pose methods, poses are treated as given, so computational cost remains unchanged.
@@
Figure 6: MMSE pose estimation improves heterogeneity reconstruction over MLE. (a) Ground-truth density maps from a synthetic dataset simulating a one-dimensional conformational transition, a standard benchmark for methods such as cryoDRGN zhong2021cryodrgn and RECOVAR gilles2025cryo . Fifty equally spaced states approximate the continuous trajectory, with colors indicating positions along the pathway (only five states are shown in the figure for clarity). (b) Example of a noisy projection image from the high-noise regime used in this study, where even ground-truth poses recover only $\sim 40\%$ of the variance for the top 30 components due to finite-sample and high-noise effects. (c) Comparison of reconstructed structures using different pose priors (ground-truth, MMSE, MLE), showing MMSE yields structures closer to the true conformation state. (d) First five principal components estimated using ground-truth poses, MMSE pose estimates, and MLE pose estimates. MMSE results closely match the ground truth in structural detail, whereas MLE reconstructions degrade beyond the first modes, becoming blurrier and less representative of the underlying variability. (e) Subspace accuracy, measured as the percentage of total variance captured (as defined in (5.3)). Ground truth achieves $\sim 40\%$, MMSE $\sim 30\%$, and MLE $\sim 25\%$. (f) Eigenvalue recovery: both methods estimate the largest true eigenvalues well, but MMSE remains accurate for smaller eigenvalues, while MLE substantially overestimates them. These results show that MMSE pose estimation consistently yields more accurate recovery of conformational variability than MLE.
We further perform a more granular evaluation of structure reconstruction at the level of individual conformational states. For this analysis, we select all projection images corresponding to a specific conformational state and embed them into the latent conformational space ($z$-space, see (5.3)) using RECOVAR, conditioned on different pose priors: ground-truth, MMSE estimates, and MLE estimates. For each subset, the reconstruction corresponding to the mean embedding serves as an estimate of that specific conformational state. These reconstructions provide a direct measure of how different orientation estimates affect the accuracy of heterogeneity structure reconstruction. Figure 6(c) demonstrates that MMSE orientation estimates produce reconstructed structures closer to the ground-truth conformation than those obtained from MLE, further highlighting the advantage of MMSE pose estimation in high-noise heterogeneity analysis. We further compared the reconstructions obtained using MMSE and MLE orientation estimates against the ground-truth conformational state. The reported local FSC resolution scores (in Å) show that the 50% quantile resolutions are 11.45 Å and 11.55 Å for MMSE and MLE, respectively, while the 90% quantiles are 18.83 Å and 21.43 Å (also see gilles2025cryo Supplementary S.J.).
@@ 6 Discussion and conclusions
The MMSE orientation estimator represents the most natural Bayes estimator in the context of cryo-EM, as it coincides with the M-step reconstruction of the EM algorithm. Nevertheless, alternative Bayes estimators associated with different loss functions could also be explored in future studies. A particularly promising direction is the direct estimation of the rotation distribution from observed images and the incorporation of this information into rotation estimation or EM-based 3D reconstruction algorithms janco2022accelerated ; xu2025misspecified . Such approaches could substantially improve rotation accuracy, especially in low-SNR regimes, as indicated by our results in Figure 2, and consequently enhance volume reconstruction quality. Further extensions may involve modeling structural uncertainties, particularly relevant for flexible or heterogeneous proteins, and addressing general pose estimation problems that jointly consider rotations and translations. Finally, leveraging the Bayesian framework to derive confidence regions for individual rotations presents an exciting opportunity to improve the interpretability and reliability of rotation estimation in cryo-EM.

### v3

@@ 4.2 Empirical results for volume reconstruction and the “Einstein from Noise” phenomenon
Figure 4: Comparison of 2D image recovery using the MMSE and the MLE rotation estimators. Iterative image recovery procedures with the MMSE estimator ($\hat{g}_{\text{MMSE}}$) and MLE estimator ($\hat{g}_{\text{MLE}}$) are defined in (4.6) and (4.7), respectively. The experiment employs a template image of Einstein and a ground truth image of Newton, both rotated in 2D over a uniform polar grid with $L=30$ samples. Each image is of size $100\times 100$ pixels, and the radial direction is discretized using $R=300$ points. The reconstructed images within the dark-orange rectangle (right panel) show superior performance with the MMSE rotation estimator, with Pearson cross-correlation (PCC) provided for each reconstructed image. The MLE and MMSE reconstructions are nearly identical at high SNR ($\sigma\to 0$), as predicted by Proposition 2.2. The SNR values used for the panels (from right to left) are $10^{-2}$, $4\times 10^{-3}$, $2\times 10^{-3}$, $7\times 10^{-4}$, and $2\times 10^{-4}$. At very low SNR ($\sigma\to\infty$), the “Einstein from Noise” effect appears, where the estimator resembles the template image of Einstein rather than the underlying truth of Newton. In intermediate SNR ranges, using the MMSE estimator in the iterative step clearly outperforms the MLE estimator.
Figure 5: Comparison of 3D structure reconstruction in cryo-ET subtomograms averaging using the MMSE and MLE rotation estimators. Iterative structure reconstruction procedures with the MMSE estimator ($\hat{g}_{\text{MMSE}}$) and MLE estimator ($\hat{g}_{\text{MLE}}$) are defined in (4.6) and (4.7), respectively. At high SNR levels, a low-resolution structure emerges due to finite grid sampling of the rotation group $\mathsf{SO}(3)$, effectively acting as a low-pass filter. The 3D reconstruction using the MMSE estimator consistently outperforms the reconstruction using the MLE estimator. For high SNR conditions (i.e., $\sigma\to 0$), both estimators yield similar 3D structures, as expected from Proposition 2.2. The SNR values used for the panels (from right to left) are $10^{-2}$, $2\times 10^{-3}$, $7\times 10^{-4}$, and $2\times 10^{-4}$, with a volume size of $32\times 32\times 32$. The boxes highlighted in purple resemble the true input structure (Ribosome-S80 Wong et al. (2014)), while the orange-highlighted boxes are more similar to the initial template (beta-galactosidase Bartesaghi et al. (2014)), illustrating the “Einstein from Noise” phenomenon. Notably, at very low SNRs, the “Einstein from Noise” effect is evident with the MLE estimator but not with the MMSE estimator.
@@
A few observations can be made from Figures 4 and 5. First, in the case of high SNR (i.e., as $\sigma\to 0$), the volume reconstruction is similar whether using the MLE estimator or the MMSE estimator. This similarity is theoretically supported by Proposition 2.2. However, as the SNR decreases, the reconstructions diverge, with the volumes reconstructed using the MMSE estimator showing a better correlation with the true volume. Second, in scenarios of extremely low SNR, where the structural signal is nearly nonexistent, the phenomenon known as “Einstein from Noise” manifests in both 2D and 3D contexts. This phenomenon pertains to the inherent model bias within the reconstruction procedure, specifically in relation to the initial templates. In such cases, the reconstructed volume exhibits structural similarities to the initial template, even though the observations do not substantiate this outcome. The generation of a structured image from entirely noisy data has attracted considerable attention, particularly during a significant scientific debate regarding the structure of an HIV molecule Mao et al. (2013b); Henderson (2013); van Heel (2013); Subramaniam (2013); Mao et al. (2013a); for a comprehensive description and statistical analysis, see Balanov et al. (2024, 2025a, 2025b, 2025c). Notably, our empirical evidence suggests that the “Einstein from Noise” phenomenon is more pronounced when adopting the MLE estimator compared to the MMSE estimator, implying that the MMSE approach is less vulnerable to the choice of the initial template. Furthermore, our experiment suggests that the advantage of using the Bayesian MMSE estimator over the MLE estimator is more significant in 3D structure reconstruction tasks compared to 2D image recovery, where the 3D setting is a problem of greater interest to researchers in structural biology.
@@ 5 Structural heterogeneity analysis
Understanding heterogeneity is central to revealing the dynamic behavior of macromolecular complexes in cryo-EM. Unlike traditional 3D classification methods that assume a small number of discrete states, modern approaches aim to capture continuous structural variability by embedding projection images into a latent conformational space. Here, we adopt the RECOVAR framework Gilles and Singer (2025), which performs heterogeneity analysis using principal component analysis (PCA) based on a regularized estimate of the conformational covariance matrix. The key idea is to infer the covariance of 3D volumes directly from noisy 2D projection images, leveraging the fact that although the volumes themselves are unobserved, their projections contain sufficient statistical information.
@@ 5.1 Fixed-pose methods
A central computational assumption in RECOVAR is that the poses $\{\phi_{i}\}$ of all particle images are known and fixed, typically obtained from a consensus refinement procedure. This “fixed-pose” assumption underlies many modern heterogeneity analysis pipelines, including 3D Classification in cryoSPARC Punjani et al. (2017), 3D Variability Analysis (3DVA) Punjani and Fleet (2021), 3DFlex Punjani and Fleet (2023), cryoDRGN Zhong et al. (2021), CryoDRGN-AI-fixed Levy et al. (2024), Opus-DSD Luo et al. (2023), and RECOVAR Gilles and Singer (2025). These methods are collectively referred to as fixed-pose methods, since they treat poses as known inputs during downstream inference. Among them, RECOVAR has recently emerged as one of the most effective approaches. According to the CryoBench evaluation Jeon et al. (2024), which benchmarks multiple methods across diverse synthetic and experimental datasets, RECOVAR consistently achieves state-of-the-art performance in both structural resolution and latent space recovery. Notably, CryoBench evaluations were performed by external users, indicating RECOVAR’s strong out-of-the-box performance and robustness to hyperparameter choices. This motivates our choice to build upon the RECOVAR framework.
@@ 5.2 The mathematical model
The forward model (Gilles and Singer, 2025, Eq. (1)) used in RECOVAR is expressed in the Fourier domain as
@@
where $\tilde{y}_{i}\in\mathbb{C}^{N^{2}}$ is the observed Fourier-transformed image (with $N$ denoting the grid size along each spatial axis), $\tilde{V}_{i}\in\mathbb{C}^{N^{3}}$ is the (unknown) 3D Fourier volume corresponding to a particular conformation, $\hat{P}(\phi_{i})$ is the tomographic projection operator from 3D to 2D after a rigid-body motion $\phi_{i}=(g_{i},t_{i})$ with $g_{i}\in\mathsf{SO}(3)$ a rotation and $t_{i}\in\mathbb{R}^{2}$ an in-plane shift, $C_{i}$ is the contrast transfer function (CTF), and $\tilde{\varepsilon}_{i}$ is additive noise.
@@
where the expectation is taken over the conformational distribution. The objective of RECOVAR is to estimate both $\mu$ and $\Sigma$ from the observed projections $\tilde{y}_{i}$} with given poses $\phi_{i}$}, using regularized least-squares minimization (Gilles and Singer, 2025, Eqns. (2)–(3)).
@@ 5.3 Numerical experiments
In our experiments, we consider a synthetic one-dimensional conformational transition, a standard benchmark for continuous heterogeneity analysis methods such as cryoDRGN Zhong et al. (2021) and RECOVAR Gilles and Singer (2025). The corresponding ground-truth density maps along the conformational coordinate are shown in Figure 6(a). To generate the projection images, the underlying conformational states were uniformly sampled along the one-dimensional transition, and molecular orientations were sampled from a uniform distribution over $\mathsf{SO}(3)$. Figure 6(b) further displays a representative noisy projection under high-noise conditions.
To avoid storing and manipulating the full high-dimensional covariance matrix $\Sigma\in\mathbb{C}^{N^{3}\times N^{3}}$, RECOVAR first estimates a low-dimensional subspace of rank $r\ll N^{3}$ in which the covariance is approximated. Following (Gilles and Singer, 2025, Appendix A.2), we estimate $\hat{\Sigma}$ via regularized least-squares, then form a rank-$d$ approximation $\hat{\Sigma}_{\mathrm{col}}$ by selecting a subset of columns using a greedy SNR-based criterion that ensures both high SNR and low inter-column correlation. An orthonormal basis $\tilde{U}\in\mathbb{C}^{N^{3}\times r}$ spanning $\hat{\Sigma}_{\mathrm{col}}$ is obtained via randomized SVD. The conformational mean $\hat{\mu}$ is also estimated at this stage. Each conformational volume is then represented as
@@
where $\hat{\mu}$ is the estimated conformational mean from the first stage, and $z_{i}\in\mathbb{C}^{r}$ are low-dimensional coordinates. In RECOVAR, these coordinates are not explicitly recovered; their covariance matrix $\hat{\Sigma}_{\tilde{U}}\in\mathbb{C}^{r\times r}$ is estimated directly via a reduced least-squares problem (Eq. (14) in Gilles and Singer (2025)), yielding a low-rank covariance approximation that characterizes the variability of $z_{i}$. The eigen-decomposition of $\hat{\Sigma}_{\tilde{U}}$ then provides the estimated principal components and eigenvalues in the reduced space. The complete procedure is detailed in (Gilles and Singer, 2025, Algorithm 1).
We perform our analyses on a controlled synthetic dataset of 30,00030{,}000 projection images ($N=128$), simulated from a one-dimensional conformational transition discretized into $50$ equally spaced states to approximate continuous heterogeneity Gilles and Singer (2025). We first evaluate the effect of orientation estimation on covariance recovery. The study is conducted in the high-noise regime of RECOVAR (Figure 6(b)), where even with ground-truth poses, the top $30$ principal components capture only about $40\%$ of the variance due to finite-sample and high-noise effects. We set $r=50$ and compare three pose sources: ground truth, MMSE estimates, and MLE estimates, keeping all other steps identical. For fixed-pose methods, poses are treated as given, so computational cost remains unchanged.
@@
Figure 6: MMSE pose estimation improves heterogeneity reconstruction over MLE. (a) Ground-truth density maps from a synthetic dataset simulating a one-dimensional conformational transition, a standard benchmark for methods such as cryoDRGN Zhong et al. (2021) and RECOVAR Gilles and Singer (2025). Fifty equally spaced states approximate the continuous trajectory, with colors indicating positions along the pathway (only five states are shown in the figure for clarity). (b) Example of a noisy projection image from the high-noise regime used in this study, where even ground-truth poses recover only $\sim 40\%$ of the variance for the top 30 components due to finite-sample and high-noise effects. (c) Comparison of reconstructed structures using different pose priors (ground-truth, MMSE, MLE), showing MMSE yields structures closer to the true conformation state. (d) First five principal components estimated using ground-truth poses, MMSE pose estimates, and MLE pose estimates. MMSE results closely match the ground truth in structural detail, whereas MLE reconstructions degrade beyond the first modes, becoming blurrier and less representative of the underlying variability. (e) Subspace accuracy, measured as the percentage of total variance captured (as defined in (5.3)). Ground truth achieves $\sim 40\%$, MMSE $\sim 30\%$, and MLE $\sim 25\%$. (f) Eigenvalue recovery: both methods estimate the largest true eigenvalues well, but MMSE remains accurate for smaller eigenvalues, while MLE substantially overestimates them. These results show that MMSE pose estimation consistently yields more accurate recovery of conformational variability than MLE.
We further perform a more granular evaluation of structure reconstruction at the level of individual conformational states. For this analysis, we select all projection images corresponding to a specific conformational state and embed them into the latent conformational space ($z$-space, see (5.3)) using RECOVAR, conditioned on different pose priors: ground-truth, MMSE estimates, and MLE estimates. For each subset, the reconstruction corresponding to the mean embedding serves as an estimate of that specific conformational state. These reconstructions provide a direct measure of how different orientation estimates affect the accuracy of heterogeneity structure reconstruction. Figure 6(c) demonstrates that MMSE orientation estimates produce reconstructed structures closer to the ground-truth conformation than those obtained from MLE, further highlighting the advantage of MMSE pose estimation in high-noise heterogeneity analysis. We further compared the reconstructions obtained using MMSE and MLE orientation estimates against the ground-truth conformational state. The reported local FSC resolution scores (in Å) show that the 50% quantile resolutions are 11.45 Å and 11.55 Å for MMSE and MLE, respectively, while the 90% quantiles are 18.83 Å and 21.43 Å (also see Gilles and Singer (2025) Supplementary S.J.).
@@ 6 Discussion and conclusions
The MMSE orientation estimator represents the most natural Bayes estimator in the context of cryo-EM, as it coincides with the M-step reconstruction of the EM algorithm. Nevertheless, alternative Bayes estimators associated with different loss functions could also be explored in future studies. A particularly promising direction is the direct estimation of the rotation distribution from observed images and the incorporation of this information into rotation estimation or EM-based 3D reconstruction algorithms Janco and Bendory (2022); Xu et al. (2025). Such approaches could substantially improve rotation accuracy, especially in low-SNR regimes, as indicated by our results in Figure 2, and consequently enhance volume reconstruction quality. Further extensions may involve modeling structural uncertainties, particularly relevant for flexible or heterogeneous proteins, and addressing general pose estimation problems that jointly consider rotations and translations. Finally, leveraging the Bayesian framework to derive confidence regions for individual rotations presents an exciting opportunity to improve the interpretability and reliability of rotation estimation in cryo-EM.